A newsletter for the project ‘Population dynamics, including epidemiology’ being part of the NWO priority program ‘Nonlinear Systems’.

March 6, 1998

Now also available on WWW

http://www.bio.vu.nl/thb/users/nls

This newsletter will appear once every three months during the course of the project. Its content will mainly consist of announcements of activities, such as seminars and workshops. Moreover, information about foreign visitors being in the Netherlands as part of the program will be given.

To make this newsletter as informative as possible we ask you to let us know about activities that are planned and the date visitors are coming and give lectures. Also addresses of persons (in the Netherlands or abroad) who might be interested in receiving this newsletter are welcome.

All correspondence should be sent to: Bob Kooi

Theoretical Biology Vrije Universiteit De Boelelaan 1087 1081 HV Amsterdam

                                                   Email: kooi@bio.vu.nl

Contents

1 Editorial 2

2 Reports 2

2.1 Report on the Second Winterschool on Population Dynamics January 7 - 11 1998, Zeist, the Netherlands . . . . . . . . . . . . . . . . . . . 2

3 Master Class Spatio-Temporal patterns 4

4 Calendar 6

Contributions to this newsletter were made by: Frans Jacobs, Hans Metz.

1 Editorial

The URL of the home-page for the NLS Population Dynamics and Epidemiology project is: http//:www.bio.vu.nl/thb/users/nls It is possible to download the Newsletter in PostScript format from this home page. For the time being the newsletter will also be send by ordinary mail.

This home-page will be updated four times a year when also the new newsletter is issued. Only the calendar is updated more frequently. Please do not hesitate to contact for updates, improvements, suggestions and so on.

2 Reports

2.1 Report on the Second Winterschool on Population Dynamics January 7 - 11 1998, Zeist, the Netherlands

by Frans Jacobs (Institute of Evolutionary and Ecological Sciences, University of Leiden)

The winterschool was organised by: Odo Diekmann (Mathematical Institute, University of Utrecht), Hans Heesterbeek (Centre for Biometry, Wageningen), Hans Metz (Institute of Evolutionary and Ecological Sciences, University of Leiden), Gerard Mulder (Institute of Evolutionary and Ecological Sciences, University of Leiden). Supported by the priority program Non-Linear Systems of the Dutch Organization for Science NWO.

About 55 young scientists from different nationalities came to Zeist in the beginning of 1998 to participate in the Second Winterschool on Population Dynamics. Inspired by the success of the first Winterschool, two years ago, this Winterschool was organized along the same lines. Five speakers were invited:

Ulf Dieckmann (IIASA, Laxenburg)
Mats Gyllenberg (Department of Mathematics, University of Turku)
Josef Hofbauer (Institute of Mathematics, University of Vienna)
Olof Leimar  (Department of Zoology, Stockholm University)
Maurice Sabelis (Department of Pure and Applied Ecology, University of Amsterdam)

who each in their five highly educational and inspiring lectures concentrated on various aspects of population- and evolutionary dynamics.

Ulf Dieckmann

Ulf Dieckmann started his lectures by mentioning the dichotomies of the notion of fitness, a central concept in evolutionary thinking. After a discussion on evolutionary game theory and its shortcomings, he set out to enunciate the theory of adaptive dynamics: a theory of phenotypic evolution that goes beyond classical game theory. Evolutionary replicator dynamics were introduced as one of the origins for the theory of adaptive dynamics. Several stochastic and deterministic models of adaptive dynamics were derived, together with their relations. Furthermore pairwise invadibility plots as tools to conclude whether or not phenotypes can coexist, were introduced. In his last session Ulf Dieckmann discussed the various time-scales (on the level of resident population dynamics, mutation processes, adaptation and branching) that play a role in evolutionary processes.

Mats Gyllenberg

Mats Gyllenberg addressed problems related to metapopulation dynamics. His first lecture dealt with the classical (unstructured) metapopulation model of Levins (1969). Also the relation between this model and the basic Susceptible-Infected-Susceptible model of epidemiology was mentioned. After this introduction attention shifted to structured metapopulation models, structured here referring to differences in patch- and local population qualities. Central in this approach of metapopulations is the so-called rescue effect: an increased immigration rate between patches decreases the extinction rate of local populations. The analogy with structured population models was discussed. Several mechanisms at the local patch level (such as patch destruction, colonization, feedback mechanisms) have to be included in metapopulation models to come to a realistic point of view. As a mathematical tool to handle two basic ingredients in metapopulation models, local population growth and colonization, the semigroup formalism was presented. In the last lectures the relation of metapopulation dynamics with biodiversity and species extinction were discussed.

Josef Hofbauer

The lectures presented by Josef Hofbauer gave a solid mathematical foundation for explaining various phenomena that may occur in ecological models. In his first lectures aspects of monotone flows and their role in competition and cooperation models were discussed. Ecological applications that were addressed dealt with mutualism and species dispersal between a number of patches. Other important topics of his lectures were persistence and permanence of ecological systems. Also heteroclinic cycles and their attracting or repelling behaviour were covered by mathematical means. This series of lectures ended with a treatise on lowdimensional Lotka-Volterra models.

Olof Leimar

The lectures of Olof Leimar combined several topics of evolutionary game theory, life-history theory and population dynamics. In his first lecture he pointed out the differences between Darwinian fitness and invasion fitness, and how they are connected by means of demographic stability. The hawk-dove game served as a device to introduce evolutionary game theory and the notion of unbeatable strategies. The relation between game theory and adaptive dynamics was discussed. Symmetric and asymmetric games were presented, as well as repeated games.

The sequential assessment game, a model of fighting behaviour, showed an application of evolutionary game theory. Invasion fitness was explained by means of the example of the problem of sex-ratio and reproductive value of Trivers and Willard. Finally, games with time-structure (in which players have the opportunity to react to each others behaviour) and reciprocity (a means to deal with variation in partner equality) were introduced as models that deal with social interactions.

Maurice Sabelis

Maurice Sabelis raised the appetite for his lectures by starting with nice slides on his research on the evolution of plant - herbivore mutualism. The dilemmas that appear on each level of a tritrophic plant - herbivore system (milker - killer dilemma for the predators, defence investment dilemma for the plant) were addressed, resulting in a clear insight into the problems that are related to the evolution of plant defence and exploitation strategies in these systems. Coevolutionary processes of patch selection strategies of predator and prey were discussed, as well as the consequences this may have for ecological stability. In the last lecture recent work on the evolutionary consequences of spatial patterns in a host - parasitoid system was presented.

Participants were provided with two readers that contained papers of the lecturers and others, dealing with the presented topics.

To get each other acquainted with one another’s work, each participant had to introduce her or his work during a short presentation. This showed that the Winterschool attracted people from all kinds of backgrounds, ranging from population biology to health research to nature conservation. During the afternoon sessions small groups were formed to study and discuss working papers that related to topics the lecturers had introduced. Each group presented the results of the discussions in public.

The five days of the Winterschool had a high density of lectures. The evenings however were outstanding occasions to wash away all the information to the deeper regions of the brains and to have conversations on other matters of life over a glass of beer or a B-52 cocktail, one of the specialties of the bartender.

3 Master Class Spatio-Temporal patterns

What follows is the final schedule for the spring semester of the Master Class Spatio-Temporal patterns. June 2–6 1998 a final Summer School ”Spatio-Temporal Patterns” Twente University, the Netherlands with five speakers. See Calendar for more details.

x means two times 45 minutes, □ means a full day. Morning classes are from 11.00-13.00; afternoon classes are from 14.00-16.00, except in week 11-16 on thursdays and fridays: On those days the schedule is 12.00-13.00 and 14.00-17.00 U = Utrecht, G = Groningen, A = Amsterdam, 1 = Monday, 2 = Tuesday, 3 = Wednesday, 4 = Thursday, 5 = Friday.

week  5  6  7  8  9  10  11  12  13  14  17  18 23

numerical mathematics: model-properties-solutions

□U4 □U4 □U4 □U4 □U4 □U4 s

(non)linear statistical analyses of processes in space and time

□U5 □U5 □U4 □U4 □U5 □U5 u

seminar in preparation of the final summer school (a)

xU3 xU3 xU3 xU3 xU3 xU3 xU3 xU3 xU3 xU3 xU3 xU3 m

ergodic theory (m)  xU1 xU1 xU1  m

empirical orthogonal functions (a)

xG1 xG1 e

visualization (m)  xG1 xG1 □G1  xG1  r

experimental analyses dynamical systems (a)

□G2 □G2 s
variational methods (m)  xU3 xU3 xU3  c
meteorology & chaos (m)  xU3 xU3 xU3 xU3 xU3 xU3  h
morphology  □A1 □A1 o
ecology (a)  □U2  o
heartfibrilation (a)  xU1 xU1  xU3 xU3 l

4 Calendar

March 10, and 13, 1998

Lectures at the EEW by Peter Turchin on NonLinear Ecological Time Series Analysis

Address: Van der Klaauw Laboratory complex Sterrewachtlaan, next to Kaiserstraat 63

(20 min walking from the railway station)

lecture theatre Sterrewachtlaan 6

(oldfashioned buildings at the left hand side of the Sterrewachtlaan, second external door, pass the foundation of the telescope to the corridor, then to the left) ‘grote collegezaal’ (in the blue building)

The lectures start at 10 h. Please register, with your name, address and e-mail address for last minute announcements, with Yvon Zitman (sbqhar@rulsfb.leidenuniv.nl; Kaiserstraat 63, 2311 GP Leiden) Information: Hans Metz, Section Theoretical Biology Institute of Evolutionary and Ecological Sciences, EEW, Rijksuniversiteit Leiden. E-mail: metz@rulsfb.leidenuniv.nl

April 1, 2, and 3, 1998

Workshop at the EEW by Peter Turchin on NonLinear Ecological Time Series Analysis

Address: Van der Klaauw Laboratory complex Sterrewachtlaan, next to Kaiserstraat 63

(20 min walking from the railway station)

colloquium room Sterrewachtlaan 5

(oldfashioned buildings at the left hand side of the Sterrewachtlaan, last external door, upstairs, door at the town moat side of the building)

The workshops start at 10 h, except for the first session of the workshop (April 1), which starts at 11 h.

Please register, with your name, address and e-mail address for last minute announcements, with Yvon Zitman (sbqhar@rulsfb.leidenuniv.nl; Kaiserstraat 63, 2311 GP Leiden)

Information: Hans Metz, Section Theoretical Biology Institute of Evolutionary and Ecological Sciences, EEW, Rijksuniversiteit Leiden. E-mail: metz@rulsfb.leidenuniv.nl

June 2–6, 1998

Summer School ”Spatio-Temporal Patterns” will be held at Twente University, the Netherlands.

This course is intended for young researchers with a strong mathematical interest, and should provide them with a good overview of various approaches to understand patterns displayed by the space-time characteristics of physical, chemical, biological, ... phenomena. There will be five main lecturers, and each of them will lecture 6 hours. Their names, the title of their lectures and a short abstract are listed below.

Right at the beginning of the course, participants will be asked to present themselves, their background and scientific interests, in a “3 minutes/1 overhead transparency” presentation. We expect 30-40 participants.

Participants and lecturers will be lodged in hotel ”de Logica”. Prices are Dfl. 50,- per day for a double room and Dfl. 65,- per day for a single room. As the number of rooms is limited, we ask people to share a room as much as possible. The organisers cannot provide any financial support (neither for travel, bed and breakfast, nor for subsistence). We therefore have kept the admission fee as low as possible: Dfl. 100,–.

Applications can be send to: M.Plekenpol@math.utwente.nl Admission is on the basis of arrival of the application.

The course is sponsored by MRI, the SWON-NWO program ”Dynamical Systems and Pattern Formation” and is organised by Odo Diekmann, Arjen Doelman, Stephan van Gils and Ronald Meester. The lecturers are:

Rick Durrett (Cornell)  Stochastic and Spatial Models
Jean-Pierre Eckman (Geneva) Some topics in the study of partial differential

equations on the infinite line

Nick Ercolani (Tucson)  tba
Satya Majumbar (Bombay)  Self-organised Criticality
Yasumasa Nishiura (Sapporro) Asymptotic and transient dynamics in pattern

formation of nonlinear dissipative systems

June 16–19, 1998

Sergio Rinaldi will visit the Netherlands (Vrije Universiteit and RIZA Lelystad). In this period a lecture of Rinaldi is planned. We will organize a mini-symposium on Population Dynamics. You are invited to suggest contributers to this mini-symposium. Information: Bob Kooi, Theoretical Biology, Vrije Universiteit De Boelelaan 1087, 1081 HV Amsterdam and Marten Scheffer, RIZA Lelystad. E-mail: kooi@bio.vu.nl

June 21–26, 1998

Fifth International Conference ‘Mathematical Population Dynamics Zakopane, Poland, June 21–26, 1998. This conference will be an interdisciplinary forum where applied mathematicians, statisticians, engineers, computer scientists, biologists, epidemiologists, biomedical scientists and ecologists can share their experiences.

Web site: http://www.cyf-kr.edu.pl/zakopane/index.html

Information: Mr. Jaroslaw Smieja, Department of Automatic Control, Silesian Technical University, Akademicka 16, 44–101 Gliwice, Poland.

E-mail: JSMIEJA@peach.ia.polsl.gliwice.pl

September 4–8, 1998

Alcala First International Conference on Mathematical Ecology, Alcala de Henares (Madrid) Spain.

Web site: http://www.ciencias.alcala.es/depmat/aicme.htm

Information: Rafael Bravo de la Parra Departamento de Matematicas Universidad de Alcala 28871 Alcala de Henares (Madrid) Spain,

E-mail: aicme@mat.alcala.es

September 20–24, 1998

On Growth and Form: Spatio-Temporal Patterning in Biology Centre for Nonlinear Systems in Biology, University of Dundee, Scotland, 20–24 September 1998.

An international meeting supported by the Society for Mathematical Biology and the European Society for Mathematical and Theoretical Biology.

The meeting will mark the occasion of the 50th anniversary of the death of D’Arcy Wentworth Thompson and the topics under discussion at the meeting will be pattern formation in developmental biology, pattern formation in reaction-diffusion systems, pattern formation in intercellular systems, pattern and form in morphometrics.

Web site: http://www.mcs.dundee.ac.uk:8080/~darcy98/

Information: Dr. M.A.J. Chaplain, Department of Mathematics, University of Dundee, Dundee DD1 4HN, Scotland.

E-mail: darcy98@mcs.dundee.ac.uk or chaplain@mcs.dundee.ac.uk

29 June–3 July, 1999

The international conference on Theory and Mathematics in Biology and Medicine, to be organised in Amsterdam June 29–July 3, 1999 has as its goal to stimulate interdisciplinary collaboration between mathematicians and the bioscientists and to act as the main forum for the exchange of recent research results and new research perspectives to the widest possible community in theoretical biology and medicine in Europe and beyond.

This is the 4th of the official congresses, taking place every 3 years under the auspices of the European Society for Mathematical and Theoretical Biology (ESMTB). The meeting is unique in that it will also be the annual conference of the Society for Mathematical Biology (SMB, North America) for 1999.

The conference will be organised by the Netherlands Society for Theoretical Biology, the oldest society for theoretical biology in the world. Through these links, this will be the first World Congress devoted to one of the fastest growing domains in science where experimental biology and medicine, biochemistry, mathematics, computational science, physics and various fields of technology come together to attack many problems of great relevance to society.

Web site: http://www.bio.vu.nl/tmbm99/